# Multispecies inhomogeneous $t$-PushTASEP from antisymmetric fusion

### v2

@@ 7. ASEP Markov matrix from transfer matrix
Using $(1-t)\frac{d}{dz}\left.\log\Lambda^{1}(z|{\bf x}={\bf 1})\right|_{z=1}=-tL$ along with (93a) and (92), one can check the stationarity condition $H_{\text{ASEP}}|\overline{\mathbb{P}}({\bf m})\rangle_{{\bf x}={\bf 1}}=0$ as desired. The (unnormalized) stationary probability is given by the matrix product formula (81b) with the homogeneous specialization $x_{1}=\cdots=x_{L}=1$.
Appendix A $S^{k,1}(z)$ from antisymmetric fusion
In this appendix we write the elements $S(z)^{{\bf e}_{a},{\bf e}_{b}}_{{\bf e}_{i},{\bf e}_{i}}$ of $S^{1,1}(z)$ in (34) simply as $S(z)^{a,b}_{i,j}$, i.e.,
$S(z)^{i,i}_{i,i}=1-tz,\quad S(z)^{i,j}_{i,j}=(1-z)t^{[i>j]}\;(i\neq j),\quad S(z)^{j,i}_{i,j}=(1-t)z^{[i<j]}\;(i\neq j).$ (94)
See Figure 3.
$z$$z$$z$$z$$i$$j$$a$$b$$i$$i$$i$$i$$i$$j$$i$$j$$i$$j$$j$$i$$S(z)^{a,b}_{i,j}$$1-tz$$(1-z)t^{[i>j]}$$(1-t)z^{[i<j]}$
Figure 3. Nonzero elements of $S^{1,1}(z)$ where $0\leq i\neq j\leq n$.
Recall that $V^{1}$ has been identified with the space of local states of the $t$-PushTASEP ${\mathsf{V}}=\bigoplus_{i=0}^{n}\mathbb{C}{\mathsf{v}}_{i}$ as in (15). We start from the basic $R$-matrix $S(z)=\sum_{a,b,i,j=0}^{n}S(z)^{a,b}_{i,j}E_{ai}\otimes E_{bj}\in\mathrm{End}({\mathsf{V}}\otimes{\mathsf{V}})$, where $E_{ij}{\mathsf{v}}_{l}=\delta_{j,l}{\mathsf{v}}_{i}$. Let $P(\mathsf{u}\otimes\mathsf{v})=\mathsf{v}\otimes\mathsf{u}$ be the transposition. From (34), one sees that the image $\mathrm{Im}PS(t^{-1})=\bigoplus_{0\leq i<j\leq n}\mathbb{C}({\mathsf{v}}_{i}\otimes{\mathsf{v}}_{j}-{\mathsf{v}}_{j}\otimes{\mathsf{v}}_{i})$ is the space of antisymmetric tensors. The Yang-Baxter equation multiplied with $P_{1,2}$ from the left reads $P_{1,2}S_{1,2}(x)S_{1,3}(y)S_{2,3}(y/x)=S_{1,3}(y/x)S_{2,3}(y)P_{1,2}S_{1,2}(x)$. The choice $x=t^{-1}$ here implies that $S_{1,3}(ty)S_{2,3}(y)\in\mathrm{End}({\mathsf{V}}\otimes{\mathsf{V}}\otimes{\mathsf{V}})$ preserves the space $\left(\mathrm{Im}P_{1,2}S_{1,2}(t^{-1})\right)\otimes{\mathsf{V}}$ $=\left(\bigoplus_{0\leq i<j\leq n}\mathbb{C}({\mathsf{v}}_{i}\otimes{\mathsf{v}}_{j}-{\mathsf{v}}_{j}\otimes{\mathsf{v}}_{i})\right)\otimes{\mathsf{V}}\subset{\mathsf{V}}\otimes{\mathsf{V}}\otimes{\mathsf{V}}$.
Consider the following operator with an overall scalar factor $d_{k}(z)^{-1}$ which is included to validate the forthcoming Theorem 9:
$d_{k}(z)^{-1}S_{1,k+1}(z)S_{2,k+1}(t^{-1}z)\cdots S_{k,k+1}(t^{-k+1}z)\in\mathrm{End}({\mathsf{V}}^{\otimes k}\otimes{\mathsf{V}})\quad(0\leq k\leq n+1),$ (95)
$d_{k}(z)=(1-t^{-1}z)(1-t^{-2}z)\cdots(1-t^{-k+1}z).$ (96)
By extending the above argument, one can show that (95) can be restricted to $\bigwedge^{k}\otimes{\mathsf{V}}$, where $\bigwedge^{k}$ is the subspace of ${\mathsf{V}}^{\otimes k}$ spanned by the degree $k$ antisymmetric tensors
$\bigwedge\nolimits^{\!k}$ $=\bigoplus_{{\bf I}\in\mathscr{T}^{k}}\mathbb{C}\,\nu_{\bf I},\qquad\nu_{\bf I}=\sum_{\sigma\in\mathfrak{S}_{k}}\mathrm{sgn}(\sigma){\mathsf{v}}_{I_{\sigma(k)}}\otimes\cdots\otimes{\mathsf{v}}_{I_{\sigma(1)}}\in{\mathsf{V}}^{\otimes k},$ (97)
where $\mathfrak{S}_{k}$ is the symmetric group of degree $k$, and $\mathrm{sgn}(\sigma)$ denotes the signature of the permutation $\sigma$. The set $\mathscr{T}^{k}$ has been defined in (16). We identify $\bigwedge^{k}$ in (97) with $V^{k}$ in (14) via
$V^{k}\ni v_{\bf i}=\nu_{\bf I}\in\bigwedge\nolimits^{\!k},\quad\mathscr{B}^{k}\ni{\bf i}={\bf I}\in\mathscr{T}^{k},$ (98)
where the bijective correspondence between the multiplicity arrays in $\mathscr{B}^{k}$ (13) and the column strict tableaux in $\mathscr{T}^{k}$ has been explained after (16).
The $R$-matrix $S(z)=S^{k,1}(z)$ in (29) is obtained as the restriction of (95) to $\bigwedge^{k}\otimes{\mathsf{V}}$ according to the above identification. This construction leads to the following formula for its matrix elements:
$S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ $=d_{k}(z)^{-1}\sum_{b_{1},\ldots,b_{k-1}\in\{0,\ldots,n\}}\sum_{\sigma\in\mathfrak{S}_{k}}\mathrm{sgn}(\sigma)S(z)^{A_{k},b}_{I_{\sigma(k)},b_{k-1}}S(t^{-1}z)^{A_{k-1},b_{k-1}}_{I_{\sigma(k-1)},b_{k-2}}\cdots S(t^{-k+1}z)^{A_{1},b_{1}}_{I_{\sigma(1)},j},$ (99)
where $b,j\in\{0,\ldots,n\}$ and ${\bf a},{\bf i}\in\mathscr{B}^{k}$ are multiplicity arrays. The arrays $(A_{1},\ldots,A_{k})$ and $(I_{1},\ldots,I_{k})\in\mathscr{T}^{k}$ are the tableau representations of ${\bf a}$ and ${\bf i}$, respectively. We note that the sign factor $\mathrm{sgn}(\sigma)$ is a simplifying feature of the current $S(z)$-gauge, in contrast to the factor $(-t)^{\mathrm{length}(\sigma)}$, which is commonly encountered in the conventional $R(z)$-gauge. To summarize, the weight is given in Figure 4.
[…]
Figure 4. The weight of the element in (99). Note that $I_{1}<I_{2}<\dots<I_{k}$ and $A_{1}<A_{2}<\dots<A_{k}$ by (16).
We now prove the main result of this section. We will use the notation $[i,j]=\{i,i+1,\dots,j-1,j\}$ for $i\leq j$.
Theorem 9.
We have
$S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ $=\delta^{{\bf a}+{\bf e}_{b}}_{{\bf i}+{\bf e}_{j}}(-1)^{a_{0}+\cdots+a_{j-1}+i_{0}+\cdots+i_{b-1}}t^{a_{j+1}+\cdots+a_{n}}(1-t^{a_{j}}z^{\delta_{b,j}})z^{[j>b]}$ (100)
$=$ $\delta^{{\bf a}+{\bf e}_{b}}_{{\bf i}+{\bf e}_{j}}(-1)^{\#\{s\in[1,k]\mid A_{s}<j\}+\#\{s\in[1,k]\mid I_{s}<b\}}\;t^{\#\{s\in[1,k]\mid A_{s}>j\}}\;(1-t^{a_{j}}z^{\delta_{b,j}})z^{[j>b]}.$ (101)
@@ Appendix A $S^{k,1}(z)$ from antisymmetric fusion
We will perform induction on $k$. For $k=1$, there is only a single term in the sum. According to (101), the answer should be
@@
Now suppose the results holds for $k-1$. That is to see that for all fixed tuples $I^{\prime}=(I_{2},\dots,I_{k}),A^{\prime}=(A_{2},\dots,A_{k})\in\mathscr{T}^{k-1}$ and elements $j^{\prime},b\in\mathscr{B}^{1}$, the weight $S(z)^{{\bf A^{\prime}},{\bf e}_{b}}_{{\bf I^{\prime}},{\bf e}_{j^{\prime}}}$ given in Figure 4 is equal to (101) with $k$ replaced by $k-1$. We now consider the $k$-weight, which is like adding one more row at the bottom to the diagram in Figure 4. For consistency with the induction hypothesis, let $j^{\prime}$ be the label attached to the vertical line between the bottom two rows.
@@
$[A_{1}>j]+\#\{s\in[2,k]\mid A_{s}>j\}=\#\{s\in[1,k]\mid A_{s}>j\},$
which matches the power in (101). The factor of $1-t^{-k+1}z$ cancels and the power of $z$ is unchanged.
@@
#$\#\{s\in[1,k]\mid\sigma(I_{s})<b\}=\#\{s\in[2,k]\mid\sigma(I_{s})<b\}+[\sigma(I_{1})<b].$
If $j<A_{1}$, then $j<A_{1},\dots,A_{k}$ and so $j=b$ since it must exit somewhere. Thus $[A_{1}<j]=[\sigma(I_{1})<b]=0$. If $A_{1}<j$ and $j=b$, then $[A_{1}<j]=[\sigma(I_{1})<b]=1$ trivially. In both these cases, #$\#\{s\in[1,k]\mid I_{s}<A_{1}\}=0$ since $\sigma(I_{1})=A_{1}$ is the smallest entry in $I$. If $A_{1}<j$ and $j\neq b$, then $j=A_{u}$ for some $u\in[2,k]$. Now, #$\#\{s\in[1,k]\mid I_{s}<A_{1}\}=[A_{1}>b]$. Therefore, the total extra sign is the parity of $[A_{1}<j]+[A_{1}<b]+[A_{1}>b]$, which is even. Thus, in all cases the signs match when $A_{1}\neq j$.
@@
$\frac{1-t^{-k+2}z}{1-t^{-k+1}z}(-1)^{\#\{s\in[2,k]\mid A_{s}<j\}+\#\{s\in[2,k]\mid\sigma(I_{1})=j,\sigma(I_{s})<b\}}\;t^{\#\{s\in[2,k]\mid A_{s}>j\}}\;z^{[j>b]}\left(1-t^{[j\in\{A_{2},\dots,A_{k}\}]}z^{[b=j]}\right).$ (102)
@@
$\frac{(1-t)\,(zt^{-k+1})^{[j>j^{\prime}]}}{1-t^{-k+1}z}(-1)^{\#\{s\in[2,k]\mid A_{s}<j^{\prime}\}+\#\{s\in[2,k]\mid\sigma(I_{1})=j^{\prime},\sigma(I_{s})<b\}+\#\{s\in[1,k]\mid I_{s}<j^{\prime}\}}\\ \times\;t^{\#\{s\in[2,k]\mid A_{s}>j^{\prime}\}}\;z^{[j^{\prime}>b]}\left(1-t^{[j^{\prime}\in\{A_{2},\dots,A_{k}\}]}z^{[b=j^{\prime}]}\right).$ (103)
Notice that there is an extra contribution to the sign in (103) because of the sign of the permutation $\sigma$ in Figure 4. We want to analyze the sum of (102) and (103). This has to be done on a case-by-case basis.
First, suppose $j=b$. Then (102) becomes
$\frac{1-t^{-k+2}z}{1-t^{-k+1}z}(-1)^{0+0}t^{k-1}z^{0}(1-t^{0}z^{1})=\frac{t^{k}(t^{k-2}-z)(1-z)}{t^{k-1}-z}.$ (104)
We have to sum (103) over all possible values of $u\in[2,k]$ such that $j^{\prime}=A_{u}$ since $j^{\prime}\neq b$. Thus, $j^{\prime}>j$ and we obtain
$\sum_{u=2}^{k}\frac{(1-t)}{1-t^{-k+1}z}(-1)^{(u-2)+0+(u-1)}t^{k-u}z^{1}(1-t^{1}z^{0})=-\frac{t^{k-1}(1-t)^{2}z}{t^{k-1}-z}\sum_{u=2}^{k}t^{k-u},$ (105)
@@
$-\frac{t^{k-1}(1-t)(1-t^{k-1})z}{t^{k-1}-z}.$ (106)
Summing (104) and (106) gives $t^{k-1}(1-tz)$, which matches (101) for $A_{1}=j=b$.
@@
If $b<j$, then $j^{\prime}$ cannot equal $j$ because $j\notin\{A_{2},\dots,A_{k},b\}$. Therefore, (102) cannot contribute. We sum (103) over all possible values of $u\in[2,k]$ such that $j^{\prime}=A_{u}$, and in addition also consider $j^{\prime}=b$. In the former case, we obtain exactly (105), and in the latter,
@@
Summing this with (105) gives, after some simplifications, $t^{k-1}(1-t)z$, which matches (101) for $A_{1}=j>b$.
The last case is when $j<b$. Notice that $A_{u}$ cannot be equal to $b$ for any $u\in[2,k]$ by conservation. Therefore $b$ is a label different from $\{A_{1},\dots,A_{k}\}$. Thus, again by conservation, $j^{\prime}$ cannot equal $j$ and so (102) does not contribute. So, we need to look at (103). Suppose $b$ is such that $A_{v}<b<A_{v+1}$ for some $v\in[k]$, where we interpret $v=k$ as saying that $b>A_{k}$. As in the situation immediately above, $j^{\prime}$ can equal $A_{u}$ for some $u\in[2,k]$ or $j^{\prime}$ can equal $b$. In the former case, we get
@@
$(-1)^{v}\frac{t^{k-1}(1-t)}{t^{k-1}-z}\left(t^{k-v}(1-t^{v-1})+z(1-t^{k-v})\right).$ (107)
@@
Summing this with (107) gives $(-1)^{v-1}t^{k-1}(1-t)$, which again matches (101) for $A_{1}=j<b$.

### v3

@@ 7. ASEP Markov matrix from transfer matrix
Using $(1-t)\frac{d}{dz}\left.\log\Lambda^{1}(z|{\bf x}={\bf 1})\right|_{z=1}=-tL$ along with (7.6a) and (7.5), one can check the stationarity condition $H_{\text{ASEP}}|\overline{\mathbb{P}}({\bf m})\rangle_{{\bf x}={\bf 1}}=0$ as desired. The (unnormalized) stationary probability is given by the matrix product formula (6.8b) with the homogeneous specialization $x_{1}=\cdots=x_{L}=1$.
Appendix A Proof of Theorem 3.1
Here we prove
$S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}=\delta^{{\bf a}+{\bf e}_{b}}_{{\bf i}+{\bf e}_{j}}(-1)^{\#\{s\in[1,k]\mid A_{s}<j\}+\#\{s\in[1,k]\mid I_{s}<b\}}\;t^{\#\{s\in[1,k]\mid A_{s}>j\}}\;(1-t^{a_{j}}z^{\delta_{b,j}})z^{[j>b]},$ (A.1)
where we have expressed (3.12) in terms of $A,I\in\mathscr{T}^{k}$ which correspond to the tableau representations of ${\bf a},{\bf i}\in\mathscr{B}^{k}$.
@@
We will perform induction on $k$. For $k=1$, there is only a single term in the sum. According to (A.1), the answer should be
@@
Now suppose the results holds for k1k-1. That is to see that for all fixed tuples $I^{\prime}=(I_{2},\dots,I_{k}),A^{\prime}=(A_{2},\dots,A_{k})\in\mathscr{T}^{k-1}$ and elements $j^{\prime},b\in\mathscr{B}^{1}$, the weight $S(z)^{{\bf A^{\prime}},{\bf e}_{b}}_{{\bf I^{\prime}},{\bf e}_{j^{\prime}}}$ given in Figure 4 is equal to (A.1) with $k$ replaced by k1k-1. We now consider the $k$-weight, which is like adding one more row at the bottom to the diagram in Figure 4. For consistency with the induction hypothesis, let $j^{\prime}$ be the label attached to the vertical line between the bottom two rows.
@@
$A_{1}>j]+\#\{s\in[2,k]\mid A_{s}>j\}=\#\{s\in[1,k]\mid A_{s}>j\},$
which matches the power in (A.1). The factor of $1-t^{-k+1}z$ cancels and the power of $z$ is unchanged.
@@
$\{s\in[1,k]\mid\sigma(I_{s})<b\}=\#\{s\in[2,k]\mid\sigma(I_{s})<b\}+[\sigma(I_{1})<b].$
If $j<A_{1}$, then $j<A_{1},\dots,A_{k}$ and so $j=b$ since it must exit somewhere. Thus $A_{1}<j]=[\sigma(I_{1})<b]=0$. If $A_{1}<j$ and $j=b$, then $A_{1}<j]=[\sigma(I_{1})<b]=1$ trivially. In both these cases, $\{s\in[1,k]\mid I_{s}<A_{1}\}=0$ since $\sigma(I_{1})=A_{1}$ is the smallest entry in $I$. If $A_{1}<j$ and $j\neq b$, then $j=A_{u}$ for some $u\in[2,k]$. Now, $\{s\in[1,k]\mid I_{s}<A_{1}\}=[A_{1}>b]$. Therefore, the total extra sign is the parity of $A_{1}<j]+[A_{1}<b]+[A_{1}>b]$, which is even. Thus, in all cases the signs match when $A_{1}\neq j$.
@@
$\frac{1-t^{-k+2}z}{1-t^{-k+1}z}(-1)^{\#\{s\in[2,k]\mid A_{s}<j\}+\#\{s\in[2,k]\mid\sigma(I_{1})=j,\sigma(I_{s})<b\}}\\ \times\;t^{\#\{s\in[2,k]\mid A_{s}>j\}}\;z^{[j>b]}\left(1-t^{[j\in\{A_{2},\dots,A_{k}\}]}z^{[b=j]}\right).$ (A.2)
@@
$\frac{(1-t)\,(zt^{-k+1})^{[j>j^{\prime}]}}{1-t^{-k+1}z}(-1)^{\#\{s\in[2,k]\mid A_{s}<j^{\prime}\}+\#\{s\in[2,k]\mid\sigma(I_{1})=j^{\prime},\sigma(I_{s})<b\}+\#\{s\in[1,k]\mid I_{s}<j^{\prime}\}}\\ \times\;t^{\#\{s\in[2,k]\mid A_{s}>j^{\prime}\}}\;z^{[j^{\prime}>b]}\left(1-t^{[j^{\prime}\in\{A_{2},\dots,A_{k}\}]}z^{[b=j^{\prime}]}\right).$ (A.3)
Notice that there is an extra contribution to the sign in (A.3) because of the sign of the permutation $\sigma$ in Figure 4. We want to analyze the sum of (A.2) and (A.3). This has to be done on a case-by-case basis.
First, suppose $j=b$. Then (A.2) becomes
$\frac{1-t^{-k+2}z}{1-t^{-k+1}z}(-1)^{0+0}t^{k-1}z^{0}(1-t^{0}z^{1})=\frac{t^{k}(t^{k-2}-z)(1-z)}{t^{k-1}-z}.$ (A.4)
We have to sum (A.3) over all possible values of $u\in[2,k]$ such that $j^{\prime}=A_{u}$ since $j^{\prime}\neq b$. Thus, $j^{\prime}>j$ and we obtain
$\sum_{u=2}^{k}\frac{(1-t)}{1-t^{-k+1}z}(-1)^{(u-2)+0+(u-1)}t^{k-u}z^{1}(1-t^{1}z^{0})=-\frac{t^{k-1}(1-t)^{2}z}{t^{k-1}-z}\sum_{u=2}^{k}t^{k-u},$ (A.5)
@@
$\frac{t^{k-1}(1-t)(1-t^{k-1})z}{t^{k-1}-z}.$ (A.6)
Summing (A.4) and (A.6) gives $t^{k-1}(1-tz)$, which matches (A.1) for $A_{1}=j=b$.
@@
If $b<j$, then $j^{\prime}$ cannot equal $j$ because $j\notin\{A_{2},\dots,A_{k},b\}$. Therefore, (A.2) cannot contribute. We sum (A.3) over all possible values of $u\in[2,k]$ such that $j^{\prime}=A_{u}$, and in addition also consider $j^{\prime}=b$. In the former case, we obtain exactly (A.5), and in the latter,
@@
Summing this with (A.5) gives, after some simplifications, $t^{k-1}(1-t)z$, which matches (A.1) for $A_{1}=j>b$.
The last case is when $j<b$. Notice that $A_{u}$ cannot be equal to $b$ for any $u\in[2,k]$ by conservation. Therefore $b$ is a label different from $\{A_{1},\dots,A_{k}\}$. Thus, again by conservation, $j^{\prime}$ cannot equal $j$ and so (A.2) does not contribute. So, we need to look at (A.3). Suppose $b$ is such that $A_{v}<b<A_{v+1}$ for some $v\in[k]$, where we interpret $v=k$ as saying that $b>A_{k}$. As in the situation immediately above, $j^{\prime}$ can equal $A_{u}$ for some $u\in[2,k]$ or $j^{\prime}$ can equal $b$. In the former case, we get
@@
$(-1)^{v}\frac{t^{k-1}(1-t)}{t^{k-1}-z}\left(t^{k-v}(1-t^{v-1})+z(1-t^{k-v})\right).$ (A.7)
@@
Summing this with (A.7) gives $(-1)^{v-1}t^{k-1}(1-t)$, which again matches (A.1) for $A_{1}=j<b$.
@@
Appendix B 3D construction of $S^{k,1}(z)$
$\mathcal{L}^{0,0}_{0,0}=\mathcal{L}^{1,1}_{1,1}=1,\quad\mathcal{L}^{1,0}_{1,0}={\rm{\bf k}},\quad\mathcal{L}^{0,1}_{0,1}=-q{\rm{\bf k}},\quad\mathcal{L}^{1,0}_{0,1}={\rm{\bf a}}^{+},\quad\mathcal{L}^{0,1}_{1,0}={\rm{\bf a}}^{-},$ (B.1a)
$\mathcal{L}^{a,b}_{i,j}=0\quad\text{if}\;\;a+b\neq i+j.$ (B.1b)
Here ${\rm{\bf k}},{\rm\mathbf{a}}^{\!+},{\rm\mathbf{a}}^{\!-}$ are $q$-oscillator operators [^9] The parameter $q$ will be set to $t^{1/2}$ in (B.10). on the Fock space $F=\bigoplus_{m\geq 0}\mathbb{Q}(q)|m\rangle$, defined by
${\rm{\bf k}}|m\rangle=q^{m}|m\rangle,\quad{\rm\mathbf{a}}^{\!+}|m\rangle=|m+1\rangle,\quad{\rm\mathbf{a}}^{\!-}|m\rangle=(1-q^{2m})|m-1\rangle.$ (B.2)
[…]
Figure 6. $\mathcal{L}=(\mathcal{L}^{a,b}_{i,j})$ as a $q$-oscillator-weighted six-vertex model. The $q$-oscillators may be regarded as acting along the third arrow perpendicular to the sheet in each vertex. In this context, $\mathcal{L}$ is referred to as a 3D $L$-operator. This figure, which will only be used in this Appendix, should not be confused with Figure 3.
For $0\leq k,l\leq n+1$, we introduce the linear map $R(z)=R^{k,l}(z)\in\mathrm{End}(V^{k}\otimes V^{l})$ by
$R(z)(v_{\bf i}\otimes v_{\bf j})$ $=\sum_{{\bf a}\in\mathscr{B}^{k},{\bf b}\in\mathscr{B}^{l}}R(z)^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}\,v_{\bf a}\otimes v_{\bf b}\qquad({\bf i}\in\mathscr{B}^{k},{\bf j}\in\mathscr{B}^{l}),$ (B.3a)
$R(z)^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}$ $=\mathrm{Tr}_{F}(z^{\bf h}\mathcal{L}^{a_{n},b_{n}}_{i_{n},j_{n}}\cdots\mathcal{L}^{a_{0},b_{0}}_{i_{0},j_{0}}).$ (B.3b)
The trace is convergent as a formal power series in $q$ and $z$. From (B.1b), $R(z)$ has the weight conservation property:
$R(z)^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}=0\;\,\text{unless}\;\,{\bf a}+{\bf b}={\bf i}+{\bf j}.$ (B.4)
The cases $k=0,n+1$ reduce to the scalar matrices as
$R^{0,l}(z)$ $=\mathrm{Tr}_{F}\left(z^{\bf h}(\mathcal{L}^{0,0}_{0,0})^{n+1-l}(\mathcal{L}^{0,1}_{0,1})^{l}\right)\mathrm{Id}=\frac{(-q)^{l}}{1-q^{l}z}\mathrm{Id},$ (B.5)
$R^{n+1,l}(z)$ $=\mathrm{Tr}_{F}\left(z^{\bf h}(\mathcal{L}^{1,1}_{1,1})^{l}(\mathcal{L}^{1,0}_{1,0})^{n+1-l}\right)\mathrm{Id}=\frac{1}{1-q^{n+1-l}z}\mathrm{Id}.$ (B.6)
The first nontrivial case is $l=1$. We express the elements ${\bf b},{\bf j}\in\mathscr{B}^{1}$ as ${\bf e}_{b},{\bf e}_{j}$ with $0\leq b,j\leq n$. Then (B.3b) is evaluated explicitly as
$(1-q^{k-1}z)(1-q^{k+1}z)R(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ $=\delta^{{\bf a}+{\bf e}_{b}}_{{\bf i}+{\bf e}_{j}}\times\begin{cases}(-q)^{1-a_{j}}(1-q^{2a_{j}+k-1}z),&j=b,\\ q^{a_{j+1}\cdots+a_{b-1}}(1-q^{2}),&j<b,\\ q^{k-1-(a_{b+1}+\cdots+a_{j-1})}(1-q^{2})z,&j>b.\end{cases}$ (B.7)
$R^{k_{1},k_{2}}_{1,2}(x)R^{k_{1},k_{3}}_{1,3}(xy)R^{k_{2},k_{3}}_{2,3}(y)=R^{k_{2},k_{3}}_{2,3}(y)R^{k_{1},k_{3}}_{1,3}(xy)R^{k_{1},k_{2}}_{1,2}(x),$ (B.8)
where $k_{1},k_{2},k_{3}\in\{0,\ldots,n+1\}$. They are equalities in $\mathrm{End}(V^{k_{1}}\otimes V^{k_{2}}\otimes V^{k_{3}})$ on which $R^{k_{i},k_{j}}_{i,j}(z)$ acts on the $i$’th and the $j$’th components as $R^{k_{i},k_{j}}(z)$ and identity elsewhere. Details can be found in [K22, Chap. 11].
B.1. Modifying $R(z)$ into $S(z)$
${\tilde{S}}(z)(v_{\bf i}\otimes v_{{\bf j}})$ $=\sum_{{\bf a}\in\mathscr{B}^{k},{\bf b}\in\mathscr{B}^{l}}{\tilde{S}}(z)^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}\,v_{\bf a}\otimes v_{{\bf b}}\qquad({\bf i}\in\mathscr{B}^{k},{\bf j}\in\mathscr{B}^{l}),$ (B.9)
${\tilde{S}}(z)^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}$ $=(-q)^{k-l+\eta}R(q^{l-k}z)^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}\left|{}_{q\rightarrow t^{1/2}}\right$., (B.10)
$\eta$ $=\eta^{{\bf a},{\bf b}}_{{\bf i},{\bf j}}=\sum_{0\leq r<s\leq n}(b_{r}a_{s}-i_{r}j_{s}).$ (B.11)
The quantity $\eta$ (B.11) is formally the same as [KMMO16, eq. (16)]. Obviously, ${\tilde{S}}(z)$ also possesses the weight conservation property as (B.4). Moreover, the Yang-Baxter equation (B.8) for $R^{k,l}(z)$ and the same argument as in the proof of [KMMO16, Prop.4] imply that ${\tilde{S}}^{k,l}(z)$ also satisfies the Yang-Baxter equation:
${\tilde{S}}^{k_{1},k_{2}}_{1,2}(x){\tilde{S}}^{k_{1},k_{3}}_{1,3}(xy){\tilde{S}}^{k_{2},k_{3}}_{2,3}(y)={\tilde{S}}^{k_{2},k_{3}}_{2,3}(y){\tilde{S}}^{k_{1},k_{3}}_{1,3}(xy){\tilde{S}}^{k_{1},k_{2}}_{1,2}(x),$ (B.12)
where $k_{1},k_{2},k_{3}\in\{0,\ldots,n+1\}$.
$S(z)(v_{\bf i}\otimes v_{{\bf e}_{j}})$ $=\sum_{{\bf a}\in\mathscr{B}^{k},{\bf e}_{b}\in\mathscr{B}^{1}}S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}\,v_{\bf a}\otimes v_{{\bf e}_{b}}\qquad({\bf i}\in\mathscr{B}^{k},{\bf e}_{j}\in\mathscr{B}^{1}),$ (B.13)
$S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ $=(1-z)(1-q^{2}z)(-q)^{k-1+\eta}R(q^{1-k}z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}\left|{}_{q\rightarrow t^{1/2}}\right$., (B.14)
$\eta$ $=\eta^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}=\sum_{s>b}a_{s}-\sum_{r<j}i_{r},$ (B.15)
where, from (B.4), we assume ${\bf a}+{\bf e}_{b}={\bf i}+{\bf e}_{j}$. This leads to the expression
$k-1+\eta=\begin{cases}a_{j}-1+2(a_{j+1}+\cdots+a_{n}),&j=b,\\ -(a_{j+1}+\cdots+a_{b-1})+2(a_{j+1}+\cdots+a_{n}),&j<b,\\ (a_{b+1}+\cdots+a_{j-1})+2(a_{j+1}+\cdots+a_{n}),&j>b.\end{cases}$ (B.16)
Therefore the matrix element (B.14) takes the form, which reproduces (3.12):
$S(z)^{{\bf a},{\bf e}_{b}}_{{\bf i},{\bf e}_{j}}$ $=\delta^{{\bf a}+{\bf e}_{b}}_{{\bf i}+{\bf e}_{j}}(-1)^{a_{0}+\cdots+a_{j-1}+i_{0}+\cdots+i_{b-1}}t^{a_{j+1}+\cdots+a_{n}}(1-t^{a_{j}}z^{\delta_{b,j}})z^{[j>b]}.$ (B.17)
It is noteworthy that (B.17) is a polynomial in both $t$ and $z$.
$S(z)^{{\bf e}_{i},{\bf e}_{i}}_{{\bf e}_{i},{\bf e}_{i}}=1-tz,\quad S(z)^{{\bf e}_{i},{\bf e}_{j}}_{{\bf e}_{i},{\bf e}_{j}}=(1-z)t^{[i>j]}\;(i\neq j),\quad S(z)^{{\bf e}_{j},{\bf e}_{i}}_{{\bf e}_{i},{\bf e}_{j}}=(1-t)z^{[i<j]}\;(i\neq j).$ (B.18)
They coincide with (3.6).